# Generalized Moving Least-Squares for Solving Vector-valued PDEs on Unknown Manifolds

### v2

@@ 4.3 Reduction from ambient dimension to intrinsic dimension
::=:= $\mathbf{T}^{\otimes}(\sum_{\ell=1}^{3}\mathbf{R}_{\ell}\mathbf{R}_{\ell})\left(\mathbf{T}^{\otimes\top}\mathbf{U}_{{\mathbf{x}_{0}}}\right).$
@@
In order to find the weights $\{\bm{w}^{k}\}_{k=1}^{K}$ as in (13) to approximate $\Delta_{B}\bm{u}\left({\mathbf{x}_{0}}\right)$, we need to further analyze the structure of the differential matrix $\mathbf{R}_{\ell}$. From a geometric view point, the left multiplication of $\mathbf{T}^{\otimes\top}$and the right multiplication of $\mathbf{T}^{\otimes}$ in the matrix $\mathbf{R}_{\ell}=\mathbf{T}^{\otimes\top}\left(\mathbf{I}_{3}\otimes\mathbf{G}_{\ell}\right)\mathbf{T}^{\otimes}$ correspond to the pointwise left projection and the pointwise right projection of the $3K\times 3K$ matrix $\mathbf{I}_{3}\otimes\mathbf{G}_{\ell}$ onto the tangent spaces of $\{{\mathbf{x}_{0,k}}\}_{k=1}^{K}$. That is, the entries in the $s^{\text{th}}$ and ($K+s$)${}^{\text{th}}$ rows of $\mathbf{R}_{\ell}$ (w.r.t. the point $\mathbf{x}_{0,s}$) and in the $r^{\text{th}}$ and ($K+r$)${}^{\text{th}}$ columns (w.r.t. the point $\mathbf{x}_{0,r}$) can be calculated by the following projection to form a $2\times 2$ block submatrix $\mathbf{\hat{R}}_{sr}^{\ell}$,
@@
where $\mathbf{\hat{R}}_{\ell}$ is row and column equivalent to $\mathbf{R}_{\ell}$, that is, $\mathbf{\hat{R}}_{\ell}$ can be obtained from $\mathbf{R}_{\ell}$ in (44) by taking appropriate row and column interchanges. Here subscript11-1- 1 is to denote the elements in the first $2$ rows of the enclosed $2K\times 2K$ matrix $\mathbf{\hat{R}}_{\ell}\mathbf{\hat{R}}_{\ell}$ corresponding to the Bochner Laplacian approximation at the base ${\mathbf{x}_{0}}$. Note that there is no need for a transformation of the weights here since only the global basis $\{\bm{t}_{1}(\mathbf{x}_{i}),\bm{t}_{2}(\mathbf{x}_{i})\}_{i=1}^{N}$ has been involved in the computation for the extrinsic formulation of the Bochner Laplacian. Following the above procedure, we can compute these weights for each base point ${\mathbf{x}_{i}}$ $(i=1,\ldots,N)$ and then appropriately arrange them into a sparse $2N$ by $2N$ Bochner Laplacian matrix.
@@ 5.2 Screened Poisson problems
where $\|\cdot\|_{2}$ is the Euclidean distance, the true solution $\mathbf{u}_{X_{M}}=(\bm{u}(\mathbf{x}_{1}),\ldots,\bm{u}(\mathbf{x}_{N}))^{\top}\in\mathbb{R}^{dN\times 1}$ is a column vector, $(\mathbf{L}_{B}\mathbf{u}_{X_{M}})_{i}\in\mathbb{R}^{d\times 1}$ corresponds to the approximate value of $\Delta_{B}\bm{u}$ at $\mathbf{x}_{i}$, $\Delta_{B}\bm{u}(\mathbf{x}_{i})\in\mathbb{R}^{d\times 1}$ is the representation in the global basis at $\mathbf{x}_{i}$, $\hat{\mathbf{u}}$ is the approximate solution, $\mathbf{T}(\mathbf{x})$ $=[\bm{t}_{1}(\mathbf{x}),$$\ldots,\bm{t}_{d}(\mathbf{x})]$ $\in\mathbb{R}^{n\times d}$ and $\mathbf{\hat{T}}(\mathbf{x})=[\bm{\hat{t}}_{1}(\mathbf{x}),\ldots,\bm{\hat{t}}_{d}(\mathbf{x})]\in\mathbb{R}^{n\times d}$ correspond to the true and the approximate global bases, respectively. Here we push both tangent vectors on (approximate) manifolds forward to tangent vectors in $\mathbb{R}^{n}$ to make the FE and IE well-defined in the Euclidean norm.
@@ 5.2.1 Hodge Laplacian on 2D sphere in $\mathbb{R}^{3}$
(a) (extr.) GMLS (b) (intr.) GMLS (c) FEEC &\&& RBF (d) time
@@ 5.2.2 Bochner Laplacian on 2D torus in $\mathbb{R}^{9}$
Figure 6: Screened Poisson problems on 2D torus in $\mathbb{R}^{9}$. Random data with 4 trials. $K=55$ nearest neighbors are used. In panels (a) and (b), shown are the average FEs of the approximate Bochner Laplacian $\mathbf{L}_{B}$ using extrinsic and intrinsic methods, respectively. In panels (c) and (d), shown are the average IEs of the solutions of the screened Poisson problem (49).
@@ 5.3 Linear time-dependent equations
where $-\pi/2\leq\theta\leq\pi/2$, $-\pi\leq\phi<\pi$, $r=3.91/3.39$, $c_{0}=0.81/3.39$, $c_{2}=7.83/3.39$ and $c_{4}=-4.39/3.39$, which is the same as the RBC in [18]. The second example is a bumpy sphere (BSP) with the parametrization:
@@ 5.4 Viscous Burgers’ equation
In the third row of Fig. 9, we verify the convergence of the proposed GMLS methods for solving the linear equation and the nonlinear Burgers’ equation. The true solution is set to be $\bm{u}=0.05\cos t\sin\theta\sin\phi\frac{\partial}{\partial\theta}$ $+0.05\cos t\sin\theta\cos\phi\frac{\partial}{\partial\phi}$ for both equations. One can see that the average SEs at time $t=0.05$ over 6 independent trials decrease at least on the order of $N^{-(l-1)/2}$ when degree $l=2,3,4,5$ as the GMLS theory predicts in (7). Here, the super-convergence can still be observed using the extrinsic method as seen in previous Section 5.2 and Section 5.3 while it can be barely seen using the intrinsic method.
@@ B.2 L Laplacian
::=:= $\mathbf{T}^{\otimes}[\sum_{\ell=1}^{3}\mathbf{R}_{\ell}\left(\mathbf{R}_{\ell}+\mathbf{M}_{\ell}\right)](\mathbf{T}^{\otimes\top}\mathbf{U}_{{\mathbf{x}_{0}}}),$
@@ B.3 Hodge Laplacian
::=:= $\mathbf{T}^{\otimes}\mathbf{J}\left(\mathbf{T}^{\otimes\top}\mathbf{U}_{{\mathbf{x}_{0}}}\right).$

### v3

@@ 4.3 Reduction from ambient dimension to intrinsic dimension
$:=$ $\mathbf{T}^{\otimes}(\sum_{\ell=1}^{3}\mathbf{R}_{\ell}\mathbf{R}_{\ell})\left(\mathbf{T}^{\otimes\top}\mathbf{U}_{{\mathbf{x}_{0}}}\right).$
@@
[…]
@@
where $\mathbf{\hat{R}}_{\ell}$ is row and column equivalent to $\mathbf{R}_{\ell}$, that is, $\mathbf{\hat{R}}_{\ell}$ can be obtained from $\mathbf{R}_{\ell}$ in (44) by taking appropriate row and column interchanges. Here subscript1-1 is to denote the elements in the first $2$ rows of the enclosed $2K\times 2K$ matrix $\mathbf{\hat{R}}_{\ell}\mathbf{\hat{R}}_{\ell}$ corresponding to the Bochner Laplacian approximation at the base $\mathbf{x}_{0}$}. Note that there is no need for a transformation of the weights here since only the global basis $\{\bm{t}_{1}(\mathbf{x}_{i}),\bm{t}_{2}(\mathbf{x}_{i})\}_{i=1}^{N}$ has been involved in the computation for the extrinsic formulation of the Bochner Laplacian. Following the above procedure, we can compute these weights for each base point $\mathbf{x}_{i}$} $(i=1,\ldots,N)$ and then appropriately arrange them into a sparse $2N$ by $2N$ Bochner Laplacian matrix.
@@ 5.2 Screened Poisson problems
where $\|\cdot\|_{2}$ is the Euclidean distance, the true solution $\mathbf{u}_{X_{M}}=(\bm{u}(\mathbf{x}_{1}),\ldots,\bm{u}(\mathbf{x}_{N}))^{\top}\in\mathbb{R}^{dN\times 1}$ is a column vector, $(\mathbf{L}_{B}\mathbf{u}_{X_{M}})_{i}\in\mathbb{R}^{d\times 1}$ corresponds to the approximate value of $\Delta_{B}\bm{u}$ at $\mathbf{x}_{i}$, $\Delta_{B}\bm{u}(\mathbf{x}_{i})\in\mathbb{R}^{d\times 1}$ is the representation in the global basis at $\mathbf{x}_{i}$, $\hat{\mathbf{u}}$ is the approximate solution, $\mathbf{T}(\mathbf{x})$ $\bm{t}_{1}(\mathbf{x}),$$\ldots,\bm{t}_{d}(\mathbf{x})]$ $\in\mathbb{R}^{n\times d}$ and $\mathbf{\hat{T}}(\mathbf{x})=[\bm{\hat{t}}_{1}(\mathbf{x}),\ldots,\bm{\hat{t}}_{d}(\mathbf{x})]\in\mathbb{R}^{n\times d}$ correspond to the true and the approximate global bases, respectively. Here we push both tangent vectors on (approximate) manifolds forward to tangent vectors in $\mathbb{R}^{n}$ to make the FE and IE well-defined in the Euclidean norm.
@@ 5.2.1 Hodge Laplacian on 2D sphere in $\mathbb{R}^{3}$
(a) (extr.) GMLS (b) (intr.) GMLS (c) FEEC &\& RBF (d) time
@@ 5.2.2 Bochner Laplacian on 2D torus in $\mathbb{R}^{9}$
Figure 6: Screened Poisson problems on 2D torus in $\mathbb{R}^{9}$. In panels (a) and (b), shown are the average FEs of the approximate Bochner Laplacian $\mathbf{L}_{B}$ using extrinsic and intrinsic methods, respectively. In panels (c) and (d), shown are the average IEs of the solutions of the screened Poisson problem (49).
@@ 5.3 Linear time-dependent equations
where $\pi/2\leq\theta\leq\pi/2$, $\pi\leq\phi<\pi$, $r=3.91/3.39$, $c_{0}=0.81/3.39$, $c_{2}=7.83/3.39$ and $c_{4}=-4.39/3.39$, which is the same as the RBC in [18]. The second example is a bumpy sphere (BSP) with the parametrization:
@@ 5.4 Viscous Burgers’ equation
In the third row of Fig. 9, we verify the convergence of the proposed GMLS methods for solving the linear equation and the nonlinear Burgers’ equation. The true solution is set to be $\bm{u}=0.05\cos t\sin\theta\sin\phi\frac{\partial}{\partial\theta}$ $0.05\cos t\sin\theta\cos\phi\frac{\partial}{\partial\phi}$ for both equations. One can see that the average SEs at time $t=0.05$ over 6 independent trials decrease at least on the order of $N^{-(l-1)/2}$ when degree $l=2,3,4,5$ as the GMLS theory predicts in (7). Here, the super-convergence can still be observed using the extrinsic method as seen in previous Section 5.2 and Section 5.3 while it can be barely seen using the intrinsic method.
@@ B.2 L Laplacian
$:=$ $\mathbf{T}^{\otimes}[\sum_{\ell=1}^{3}\mathbf{R}_{\ell}\left(\mathbf{R}_{\ell}+\mathbf{M}_{\ell}\right)](\mathbf{T}^{\otimes\top}\mathbf{U}_{{\mathbf{x}_{0}}}),$
@@ B.3 Hodge Laplacian
$:=$ $\mathbf{T}^{\otimes}\mathbf{J}\left(\mathbf{T}^{\otimes\top}\mathbf{U}_{{\mathbf{x}_{0}}}\right).$